# Near-Discovery SOAR Photometry of the Third Interstellar Object: 3I/ATLAS

### v1

@@ Abstract
Interstellar Objects – Comets – Asteroids – Photometry
@@ 1 Introduction
Post-discovery observations of 3I/ATLAS revealed that it was active, reddened, and had a mostly flat light curve. Jewitt & Luu (2025) reported observations with the Nordic Optical Telescope on UT 2025 July 2 that showed that 3I/ATLAS was active. Alarcon et al. (2025) also reported observations on the night of UT 2025 July 2 from the Two-meter Twin Telescope (TTT3) at Teide Observatory that confirmed the presence of cometary activity. As additional evidence of sublimative driven mass-loss, Figure 3 in Seligman et al. (2025) showed cometary activity in deep stacks of images from UT 2025 July 2 and 4 from the ESO’s Very Large Telescope (VLT) and the Canada France Hawaii Telescope (CFHT).
@@
Recent photometric observations indicate that 3I/ATLAS has minimal brightness variation and color consistent with the first two interstellar objects. de la Fuente Marcos et al. (2025) reported the potential identification of a $P_{\rm rot}=16.79$ h rotational period from time-series photometry obtained with the OSIRIS camera spectrograph at the 10.4-meter Gran Telescopio Canarias and the Two-meter Twin Telescope. Santana-Ros et al. (2025) similarly reported a nucleus spin rate with rotational period of $P_{\rm rot}=16.16\pm 0.01$ h with a light curve amplitude of $\Delta m=0.3$ mag based on spectral and photometric observations with the SALT telescope and Nordic Optical Telescope. Seligman et al. (2025) reported additional photometric observations with ATLAS, the Faulkes Telescope North and South, LCO, Telescope Joan Oro, and TRAPPIST North and South of 3I/ATLAS. They found a light curve that was mostly flat (within the uncertainties stemming from combining data from various telescopes and reduction techniques) from UT 2025 July 2–5. They reported colors that were red similar to D-type asteroids, 1I/‘Oumuamua, and 2I/Borisov. Photometric data obtained by Beniyama (2025) on UT 2025 July 15 also indicate a red surface color and no significant light curve variability. Using photometric data from the Hubble Space Telescope (HST), Jewitt et al. (2025) confirmed cometary activity and provided an upper limit to the nuclear radius of $r_{n}<2.4$ km based on the surface brightness profile. There is an absence of observational data reported in literature thus far in the window of UT 2025 July 9 and 10. This work fills this gap and adds to an expanding composite light curve with photometric measurements of 3I/ATLAS.
@@ 2.3 Photometry
Figure 6: Light curve of 3I/ATLAS including $r^{\prime}$-band data from the SOAR telescope observations from UT 2025 July 3, 9, and 10. SOAR data are represented as octagons. The ‘X’ markers indicate points with field star contamination and ‘+’ markers indicate points with radial brightness profile contamination.
@@ 2.3.1 Aperture Experiments
Figure 8: Composite light curve of 3I/ATLAS (adapted from (Seligman et al., 2025)) including Rc-band data from the TRAPPIST Telescopes (-North and -South), o-band from the ATLAS Telescopes (Hawai’i, Sutherland, and Canary Islands), and $r^{\prime}$-band data from the LCO Telescopes, Faulkes Telescopes (-North and -South), Telescope Joan Oró, and SOAR Telescope observations from UT 2025 July 3, 9, and 10. Uncontaminated SOAR data are represented as purple octagons.
@@ 2.3.2 Stellar Contamination Search
Finally, we performed a search for stellar contamination and rejected photometric points that were significantly contaminated. In Figure 6 we show the resulting photometric data points, including points that are affected by contamination, from these three nights of SOAR data. Images in which 3I/ATLAS was contaminated were identified by both (i) visual inspection and (ii) assessment of the radial brightness profile compared to the background. Using those methods, photometric data points from UT 2025 July 3 did not show significant contamination.
@@ 2.4 Image Stacking
In Figure 7 we show the stacks of all noncontaminated images. We show stacks from UT 2025 July 09 and UT 2025 July 10 aligned on both 3I/ATLAS and field stars. There is no apparent contamination, which serves as a validation that our uncontaminated subset is correctly identified as such. We also see indications of a coma that is slightly extended approximately to the West, consistent with other published reports.
@@ 3 Composite Light Curve
It is evident from Figure 6 that there was significant scatter in the magnitude of the brightness variations in 3I/ATLAS over UT 2025 July 9 and 10. Moreover, the large brightness increases tend to cluster in time. For example, see the clusters of seemingly sporadic brightness increases at UT July 9 5:53 AM and at UT July 10 6:09 AM. This alone provides tentative evidence that the seemingly sporadic brightness increases were produced by background stellar contamination. As described in the previous section the visual inspection and PSF comparison revealed the presence of significant contamination in a large subset of our data. Therefore, each photometric measurement which was contaminated is indicated in Figure 6
In Figure 8 we show the extant photometry we obtained with each frame without the contaminated points. It is clear that the brightness increases are entirely caused by stellar contamination. It should be noted that each data point shown in Figure 8 — including these SOAR photometric data — have not been corrected for changing observing geometry. Given the distances from the first and last set of observations there is an expected brightening of $\sim 0.3$ mag. The mean brightness of 3I/ATLAS on UT 2025 July 3, 9 and 10 from these data points is 18.14, 17.55, and 17.54, respectively.
@@ 4 Discussion
It appears that these SOAR observations captured a flat light curve on UT 2025 July 3, consistent with composite light curve measurements reported to date. This implies tentatively that the production of dust — and perhaps volatiles although that would be inferred assuming some correlation of the dust and gas outburst— has been somewhat consistent from discovery through the dates when these observations were taken.
It has been noted by de la Fuente Marcos et al. (2025) and Santana-Ros et al. (2025) that the light curve as it develops is presenting a periodic signal. However, due to the relatively small temporal baseline of the new data presented here and the sparseness of it, there are many different periodic signals that could fit these data. Therefore, we do not perform a comprehensive search for periodic signals of activity in the composite light curve. This would be beneficial in future work once more data has been obtained and ideally reduced in a homogeneous manner.
@@ 5 Acknowledgments
We thank Ted Leandro De Almeida for assistance with obtaining data with SOAR. We thank Adina Feinstein and Meir Schochet for useful conservations.
@@ 6 Data and Software Availability
Software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), CALVIACat (Kelley & Lister, 2024), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020).

### v2

@@ Abstract
comets: general – techniques: photometric – minor planets, asteroids: general – telescopes
@@ 1 Introduction
Post-discovery observations of 3I/ATLAS revealed that it was active, reddened, and had a mostly flat light curve. Jewitt & Luu (2025b) reported observations with the Nordic Optical Telescope on UT 2025 July 2 that showed that 3I/ATLAS was active. Alarcon et al. (2025) also reported observations on the night of UT 2025 July 2 from the Two-meter Twin Telescope (TTT3) at Teide Observatory that confirmed the presence of cometary activity. As additional evidence of sublimative driven mass-loss, Figure 3 in Seligman et al. (2025) showed cometary activity in deep stacks of images from UT 2025 July 2 and 4 from the ESO’s Very Large Telescope (VLT) and the Canada France Hawaii Telescope (CFHT).
@@
Recent photometric observations indicate that 3I/ATLAS has minimal brightness variation and color consistent with the first two interstellar objects. de la Fuente Marcos et al. (2025) reported the potential identification of a $P_{\rm rot}=16.79$ h rotational period from time-series photometry obtained with the OSIRIS camera spectrograph at the 10.4-meter Gran Telescopio Canarias and the Two-meter Twin Telescope. Santana-Ros et al. (2025) similarly reported a nucleus spin rate with rotational period of $P_{\rm rot}=16.16\pm 0.01$ h with a light curve amplitude of $\Delta m=0.3$ mag based on spectral and photometric observations with the SALT telescope and Nordic Optical Telescope. Seligman et al. (2025) reported additional photometric observations with ATLAS, the Faulkes Telescope North and South, LCO, Telescope Joan Oro, and TRAPPIST North and South of 3I/ATLAS. They found a light curve that was mostly flat (within the uncertainties stemming from combining data from various telescopes and reduction techniques) from UT 2025 July 2–5. They reported colors that were red similar to D-type asteroids, 1I/‘Oumuamua, and 2I/Borisov. Photometric data obtained by Beniyama (2025) on UT 2025 July 15 also indicate a red surface color and no significant light curve variability. Using photometric data from the Hubble Space Telescope (HST), Jewitt et al. (2025) confirmed cometary activity and provided an upper limit to the nuclear radius of $r_{n}<2.4$ km based on the surface brightness profile. There is an absence of observational data reported in literature thus far in the window of UT 2025 July 9 and 10. This work fills this gap and adds to an expanding composite light curve with photometric measurements of 3I/ATLAS. Additional observations of 3I/ATLAS have been reported contemporaneously with those reported here (Salazar Manzano et al., 2025; Gray et al., 2025; Tonry et al., 2025; Ye et al., 2025; Hutsemékers et al., 2025; Coulson et al., 2025; Hoogendam et al., 2025; Jewitt & Luu, 2025a), including theoretical estimates (Guo et al., 2025; Pérez-Couto et al., 2025; Grant & Jones, 2025; Keto & Loeb, 2025).
@@ 2.3 Photometry
Figure 6: Light curve of 3I/ATLAS including $r^{\prime}$-band data from the SOAR telescope observations from UT 2025 July 3, 9, and 10. SOAR data are represented as octagons. The ‘X’ markers indicate points with field star contamination and ‘+’ markers indicate points with radial brightness profile contamination. The open circle markers indicate data that was not identified to be field star or profile contamination, but exhibited unusual brightness variation.
@@ 2.3.1 Aperture Experiments
Figure 8: Composite light curve of 3I/ATLAS (adapted from Seligman et al. (2025)) including Rc-band data from the TRAPPIST Telescopes (-North and -South), o-band from the ATLAS Telescopes (Hawai’i, Sutherland, and Canary Islands), and $r^{\prime}$-band data from the LCO Telescopes, Faulkes Telescopes (-North and -South), Telescope Joan Oró, and SOAR Telescope observations from UT 2025 July 3, 9, and 10. Uncontaminated SOAR data are represented as purple octagons. The dotted horizontal lines are lines of constant apparent magnitude given a constant absolute magnitude (with arbitrary zero points) and geocentric distances of 3I/ATLAS (queried from JPL Horizons).
@@ 2.3.2 Stellar Contamination Search
Finally, we performed a search for stellar contamination and rejected photometric points that were significantly contaminated. In Figure 6 we show the resulting photometric data points, including points that are affected by contamination, from these three nights of SOAR data. Images in which 3I/ATLAS was contaminated were identified by both (i) visual inspection and (ii) assessment of the radial brightness profile compared to the background. On UT 2025 July 10 at 6:00-6:15, $\sim$ 3 data points not flagged by previous contamination detection methods exhibited significant brightness variation. Upon visual inspection, we conclude that this is a result of overall brightening of the field as a result of an artifact of image reduction. These partially contaminated points are marked as open circles in Figure 6 and are removed from Figure 8. Using those methods, photometric data points from UT 2025 July 3 did not show significant contamination.
@@ 2.4 Image Stacking
In Figure 7 we show the stacks of all noncontaminated images. We show stacks from UT 2025 July 9 and UT 2025 July 10 aligned on both 3I/ATLAS and field stars. There is no apparent contamination, which serves as a validation that our uncontaminated subset is correctly identified as such. We also see indications of a coma that is slightly extended approximately to the West, consistent with other published reports from the same time period as these observations (Seligman et al., 2025; Opitom et al., 2025; Chandler et al., 2025; Cordiner et al., 2025; Jewitt & Luu, 2025a). The coma’s westward extension indicates it was in the solar direction at the time of these SOAR observations, although we note that later observations indicate that the tail has since switched to the antisolar direction (Tonry et al., 2025; Jewitt & Luu, 2025a; Hoogendam et al., 2025; Gray et al., 2025). Coma morphologies in the solar direction have been observed in Solar System comets including 19P/Borrelly which was reported to a have a strong sunward facing jet aligned with the spin axis of its nucleus (Farnham & Cochran, 2002; Farnham, 2009).
@@ 3 Composite Light Curve
It is evident from Figure 6 that there was significant scatter in the magnitude of the brightness variations in 3I/ATLAS over UT 2025 July 9 and 10. Moreover, the large brightness increases tend to cluster in time. For example, see the clusters of seemingly sporadic brightness increases at UT July 9 5:53 AM and at UT July 10 6:09 AM. This alone provides tentative evidence that the seemingly sporadic brightness increases were produced by background stellar contamination. As described in the previous section the visual inspection and PSF comparison revealed the presence of significant contamination in a large subset of our data. Therefore, each photometric measurement which was contaminated is indicated in Figure 6. Observational challenges considered in this work may have affected other near discovery photometric measurements. Future work to consolidate these data sets might provide a more homogeneous data set that corrects for these challenges and contextualizes early photometric data of 3I/ATLAS.
In Figure 8 we show the extant photometry we obtained with each frame without the contaminated points. It is clear that the brightness increases are entirely caused by stellar contamination. It should be noted that each data point shown in Figure 8 — including these SOAR photometric data — have not been corrected for changing observing geometry, including heliocentric and geocentric distances. Given the distances from the first and last set of observations there is an expected brightening of $\sim 0.3$ mag. The variation between each data set seen in the composite light curve in Figure 8 is due to differing photometric measurement methods (PSF and aperture photometry), observing geometry, zero point selection, selected aperture size, and filter. Table 2 in Seligman et al. (2025) details the different observational and photometric parameters for the TRAPPIST, LCO, ATLAS, and TJO data. For emphasis, these photometric data points are not adjusted to a common filter. This contributes to the brightness variation between different filters seen in Figure 8, specifically the SDSS $r^{\prime}$ and TRAPPIST Rc data taken contemporaneously on UT 2025 July 3. Given these variations in reduction techniques and filter used, we caution the interpretation of the dotted lines of constant apparent magnitude plotted in Figure 8. We add these lines simply to offer a more quantitative assessment of how the brightness measurements of 3I/ATLAS may have evolved in consideration of geocentric distances (queried from JPL Horizons). The mean brightness of 3I/ATLAS on UT 2025 July 3, 9 and 10 from these data points is 18.14, 17.55, and 17.54, respectively.
@@ 4 Discussion
These SOAR observations captured a flat light curve on UT 2025 July 3, consistent with composite light curve measurements reported to date. This implies tentatively that the production of dust — and perhaps volatiles although that would be inferred assuming some correlation of the dust and gas outburst— has been somewhat consistent from discovery through the dates when these observations were taken.
It has been noted by de la Fuente Marcos et al. (2025) and Santana-Ros et al. (2025) that the light curve as it develops is presenting a periodic signal. However, due to the relatively small temporal baseline of the new data presented here and the sparseness of it, it would be impossible to differentiate between any periodic signals with these data if they were present. Furthermore, the high levels of sustained activity from the coma likely suppress any noticeable photometric variation caused by the rotation of the nucleus. We explicitly do not speculate on the validity of the periodic signals identified by other authors. Therefore, we do not perform a comprehensive search for periodic signals of activity in the composite light curve. This would be beneficial in future work once more data has been obtained and ideally reduced in a homogeneous manner.
@@ 5 Acknowledgments
We thank the anonymous reviewer for their insightful comments and suggestions that strengthened the scientific content of this manuscript.
@@
We thank Ted Leandro De Almeida for assistance with obtaining data with SOAR. We thank Adina Feinstein, Meir Schochet, and Dennis Bodewits for useful conservations.
@@ 6 Data and Software Availability
Software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), CALVIACat (Kelley & Lister, 2024), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), Astrometry.net (Lang et al., 2010).